OpenStudy (anonymous):

Please! I need help finding information about three graphs? Such as length of axis, eccentricity, and length of focal chord

OpenStudy (anonymous):

I have equations for the graphs if you could help @ganeshie8

OpenStudy (anonymous):

I know some things such as the center, foci, vertices and such

ganeshie8 (ganeshie8):

conics ! il have to revise, but i can give a hand :) do u have specific question ?

OpenStudy (anonymous):

Yes! (x-5)^2/4 + (y+3)^2/16=1. I need to find the eccentricity, lengths of major and minor axis

ganeshie8 (ganeshie8):

use the chart for everything u wanto find. for eccentricity use below :- \(\large e = \frac{c}{a}\)

OpenStudy (anonymous):

use what?

ganeshie8 (ganeshie8):

use the attached chart u need to use the formulas from second column

OpenStudy (anonymous):

so length of major axis =32?

OpenStudy (anonymous):

length of minor = 8?

ganeshie8 (ganeshie8):

No. first convert the equation to one of the two standard forms.

ganeshie8 (ganeshie8):

\(\large \frac{ (x-5)^2}{4} + \frac{(y+3)^2}{16}=1\)

ganeshie8 (ganeshie8):

\(\large \frac{ (x-5)^2}{2^2} + \frac{(y+3)^2}{4^2}=1\)

ganeshie8 (ganeshie8):

a = 4 b = 2

OpenStudy (anonymous):

so major=8 and minor=4?

ganeshie8 (ganeshie8):

Length of major axis = 2a = 2*4 = 8 Length of minor axis = 2b = 2*2 = 4

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

What about eccentricity?

ganeshie8 (ganeshie8):

eccentricity = c/a u need to find c first

ganeshie8 (ganeshie8):

use below to find c :- c^2 = a^2-b^2

OpenStudy (anonymous):

c^2=20

ganeshie8 (ganeshie8):

c^2 = a^2-b^2 = 4^2 - 2^2 = 16 - 4 = 12

ganeshie8 (ganeshie8):

c = sqrt(12) = sqrt(4*3) = 2sqrt(3) eceentricity = c/a = ?

OpenStudy (anonymous):

2sqrt3/4

ganeshie8 (ganeshie8):

yes, u may simplify a bit sqrt(2)/2

ganeshie8 (ganeshie8):

* sqrt(3)/2

OpenStudy (anonymous):

Oh ok, thanks!

ganeshie8 (ganeshie8):

np :)